OpenStudy (anonymous):

The ordered pairs below represent a linear relation between x and y. (-3,1),(-2,3),(-1,5),(0,7),(1,9),(2,11). Which of the following is a correct statement about this relation? A. The slope of the line represented by this table is 2 and the y-intercept is 7. B. The slope of the line represented by this table is 2 and the y-intercept is -4. C. The slope of the line represented by this table is -2 and the y-intercept is -4. D. The slope of the line represented by this table is -2 and the y-intercept is 7.

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

im not too sure. since there are so many ordered pairs, how would i find one slope of the line and one y-intercept? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

to find the slope, pick any pair of points you want and use the slope formula

jimthompson5910 (jim_thompson5910):

the y-intercept is the point of the form (0,y)

jimthompson5910 (jim_thompson5910):

ie it's the point where the line crosses the y axis

OpenStudy (anonymous):

okay well i picked (1,9) and got y=1x+9. is that correct? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no that's not correct

OpenStudy (anonymous):

what did i do wrong?

jimthompson5910 (jim_thompson5910):

pick any two points, then use the slope formula to find the slope

OpenStudy (anonymous):

if i went with (1,9) and (2,11) would I add the two slopes of y=x+9 and y=2x+11? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

those are equations, not slopes

jimthompson5910 (jim_thompson5910):

the slope formula is m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

you use this formula to get a numeric value (not an equation)

OpenStudy (anonymous):

oh okay, so using points (1,9) and (2,11) I got 2/1 for the slope @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good, or just 2

jimthompson5910 (jim_thompson5910):

so the equation looks like y = 2x + b where b is the y-intercept

OpenStudy (anonymous):

how do i figure out the y-intercept?

jimthompson5910 (jim_thompson5910):

look for the point where the x coordinate is 0

jimthompson5910 (jim_thompson5910):

or...you can plug in any point in for x and y then solve for b

OpenStudy (anonymous):

if i wanted to solve it by plugging in an x and y could i pick any ordered pair or does it hae to be one of the ordered pairs ive already picked?

jimthompson5910 (jim_thompson5910):

it can be any ordered pair since they all lie on the same line

OpenStudy (anonymous):

okay i used (1,9) and i got: 7=b making it y=2x+7? @jim_thompson5910

OpenStudy (anonymous):

oh, so it would be A correct? :D

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

woohoo! thank you @jim_thompson5910 you're so awesome!

jimthompson5910 (jim_thompson5910):

you're welcome